OpenStudy (anonymous):

What are the possible rational zeros of f(x) = 2x3 – 4x2 – 7x + 14 ?

OpenStudy (anonymous):

\[\frac{14}{2}=\frac{\pm1,\pm2,\pm7,\pm14}{\pm1,\pm2}=\pm1,\pm\frac {1}{2},\pm \frac {7}{2}{,\pm14}\]